40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Aug 5 11:04:07 2017, gautam, Update, SUS, MC1 glitches return MC1_glitch_Aug42017.pngMC1_glitch_zoomed.png
    Reply  Wed Aug 9 15:15:47 2017, gautam, Update, SUS, MC1 glitches return MC1_glitch.png
       Reply  Thu Aug 10 21:01:43 2017, gautam, Update, SUS, MC1 glitches debugging MC_QUAD_10AUG2017.jpgMCR_10AUG2017.jpg
          Reply  Fri Aug 11 12:32:46 2017, gautam, Update, SUS, MC1 glitches debugging MC1_freeswinging.pdfMC1_inmatrix.pngMC1_sensors.png
             Reply  Fri Aug 11 17:36:47 2017, gautam, Update, SUS, MC1 <--> MC3 MC1_glitch_watchdog_shutdown.png
                Reply  Sat Aug 12 14:09:36 2017, gautam, Update, SUS, Glitches stay on MC1 MC1_glitch.pngMC_12hr_trend.pngMC1_glitches_intermittent.png
                   Reply  Sat Aug 12 22:35:22 2017, rana, Update, SUS, Glitches stay on MC1 
                      Reply  Sun Aug 13 01:35:09 2017, Koji, Update, SUS, Glitches stay on MC1 
                         Reply  Mon Aug 14 20:01:38 2017, gautam, Update, SUS, Glitches stay on MC1 
                Reply  Wed Aug 16 19:50:17 2017, gautam, Update, SUS, MC1 <--> MC3 switched back 
                   Reply  Thu Aug 17 11:17:49 2017, gautam, Update, SUS, MC1 <--> MC3 switched back MC1_misaligned.pngMC1_glitch.png
                      Reply  Thu Aug 17 17:33:01 2017, gautam, Update, SUS, MC1 <--> MC3 switched back 
Message ID: 13225     Entry time: Thu Aug 17 11:17:49 2017     In reply to: 13220     Reply to this: 13226
Author: gautam 
Type: Update 
Category: SUS 
Subject: MC1 <--> MC3 switched back 

Seems like this modification didn't really work. There were several large MC1 glitches, and one of them misaligned MC1 so much that the IMC didn't relock for the last ~6 hours. I re-aligned MC1 manually, and now it is locked fine.

Quote:

Now that all the CDS overview lights are green, I decided to switch back the coil driver outputs to their original state so that the MC optics could be damped and the IMC relocked. I also restored the static PIT/YAW bias values to their original values.

MC1 has been quiet over the last couple of days, lets see how it behaves in the next few days. In all the glitches I have observed, if the IMC is locked and WFS loops are enabled, the loops are able to correct for the DC misalignment caused by the glitch. But the mcwfs off script is currently set up in such a way that the output history is cleared between IMC locks. I made two copies of the mcwfson/mcwfsoff scripts, called mcwfsunhold/mcwfshold respectively. They live in /opt/rtcds/caltech/c1/scripts/MC/WFS. I've also modified the autolocker script to call these modified scripts, such that when the IMC loses lock, the WFS servo outputs are held, while the input is turned off. The hope is that in this configuration, the autolocker can catch a lock even if there is a glitch on MC1.

I haven't tried locking the arms yet, but I think other IFO work discussed at the meeting (like arm loss estimation / cavity scans etc) can proceed.

 

 

Attachment 1: MC1_misaligned.png  22 kB  | Hide | Hide all
MC1_misaligned.png
Attachment 2: MC1_glitch.png  43 kB  | Hide | Hide all
MC1_glitch.png
ELOG V3.1.3-